OpenStudy (anonymous):

A(1,3) and C(2,5) are two points. Find the coordinates of the point b, if c is the midpoint of [AB]

OpenStudy (nirmalnema):

A(1,3) and C(2,5) let the co ordinates of b be(x,y) (1+x)/2=2 x=3 (3+y)/2=5 y=7

OpenStudy (nirmalnema):

@sophia do u understand

OpenStudy (ahsome):

If you see, every time the x value increases by 1, the y increases by 2 Since you know that C is between A & B, and you know the difference is 1 between A & C, you know that the difference between A & B is 2 So add the x value of A with 2 x+2 1+2=3 You also know the conversion rate is x+2=y So substitute the y value x+2=y 3+2=5 Graph and check :)

OpenStudy (ahsome):

This is assuming the graph is a linear equation